### v1

# Perfect Secret Key Generation for a class of Hypergraphical Sources

Abstract
Nitinawarat and Narayan proposed a perfect secret key generation scheme for the so-called pairwise independent network (PIN) model by exploiting the combinatorial properties of the underlying graph, namely the spanning tree packing rate. This work considers a generalization of the PIN model where the underlying graph is replaced with a hypergraph, and makes progress towards designing similar perfect secret key generation schemes by exploiting the combinatorial properties of the hypergraph.
Our contributions are two-fold. We first provide a capacity achieving scheme for a complete $t$-uniform hypergraph on $m$ vertices by leveraging a packing of the complete $t$-uniform hypergraphs by what we refer to as star hypergraphs, and designing a scheme that gives $\binom{m-2}{t-2}$ bits of perfect secret key per star graph. Our second contribution is a 2-bit perfect secret key generation scheme for 3-uniform star hypergraphs whose projections are cycles. This scheme is then extended to a perfect secret key generation scheme for generic 3-uniform hypergraphs by exploiting star graph packing of 3-uniform hypergraphs and Hamiltonian packings of graphs. The scheme is then shown to be capacity achieving for certain classes of hypergraphs.
footnotetext: M. Mukherjee is with the Indraprastha Institute of Information Technology Delhi, Delhi, India. Email: manuj@iiitd.ac.in.
S. Chatterjee is with the Tata Institute of Fundamental Research, Mumbai, India. Email: chatsagnik@gmail.com.
A. Sethi is with the Indian Institute of Science, Bangalore, India. Email: alhadsethi@iisc.ac.in.
I Introduction
The multiparty secret key generation problem considers $m$ parties that need to agree upon a group secret key. To facilitate secret key generation, the parties are given access to a discrete memoryless multiterminal source (DMMS), i.e., i.i.d. copies of distinct but correlated random variables, and are allowed to communicate interactively over a noiseless but public broadcast channel. The secret key generated therefore, needs to be secured from any passive eavesdropper [^1] An eavesdropper is said to be passive if it can only listen to the messages but not tamper with them. listening to the communication sent over the channel. Initially posed for $m=2$ parties independently by [Maurer93, AC93], this problem was later extended to an arbitrary number of parties by the seminal work of Csiszár and Narayan [CN04].
The key quantity of interest in these works is the secret key capacity, i.e., the maximum rate of a secret key that can be generated, a full characterization of which appears in [CN04]. An alternate characterization of the secret key capacity was given by Chan and Zheng [CZ10] who showed that the secret key capacity equals a multivariate analogue of the mutual information [chan15], which we shall refer to as multipartite information.
Subsequent works have looked into other aspects of the multiparty secret key generation problem, such as communication rate needed to achieve secret key capacity [tyagi, mukherjee, ccde, rskmax, hypupper, pinedmond], capacity with silent parties [AG, mukherjee, zhang, so], ‘single shot’ secret key capacity [tw], and perfect secret key generation [nitinrawat, chanlin, ccde]. In this work, we focus on the aspect of perfect secret key generation.
A perfect secret key satisfies the more stringent condition of being completely independent of the communication, as opposed to the usual notion of a strong secret key where the key is allowed to be ‘almost independent’ of the communication – see Section II-A for a precise definition. While a general characterisation remains unknown, perfect secret key capacity is known to match the usual (strong) secret key capacity for certain classes of sources such as the pairwise independent network (PIN) model [pinorig].
The PIN model is defined using an underlying graph whose vertex set equals the set of parties. Associated with every edge are i.i.d. sequences of independent Be(1/2)(1/2) random variables, and each party has access to the random variables associated with the edges incident with it. Nitanawarat and Narayan [nitinrawat] gave an elegant perfect secret key generation scheme for the PIN model achieving capacity. The scheme involves packing multiple copies of the underlying graph with spanning trees, and generating a single bit of perfect secret key from every spanning tree. This scheme therefore achieves a perfect secret key of rate equal to the spanning tree packing rate of the graph, which is shown to equal the (strong) secret key capacity by exploiting standard results on spanning tree packing such as Tutte’s Theorem [diestel, Theorem 2.4.1].
A natural generalization of the PIN model is the so-called hypergraphical source [ccde] where the underlying graph is allowed to be a hypergraph. One envisages that a Nitinawarat-Narayan type scheme exploiting the combinatorial properties of the underlying hypergraph should be able to generate capacity achieving perfect secret keys. In this work, we make progress towards realizing this goal.
Our contributions in this work are two-fold. Firstly, we provide an explicit secret key achieving scheme for the complete $t$-uniform hypergraph by exploiting its combinatorial properties. We begin by developing a key generation protocol that extracts $\binom{m-2}{t-2}$ bits of perfect secret key from star hypergraphs whose projected hypergraph is complete – see Section III for formal definitions. Therefore, any packing of the complete $t$-uniform hypergraph by such star hypergraphs leads to a perfect secret key of rate equal to the packing rate times $\binom{m-2}{t-2}$. The scheme is then shown to match the secret key capacity by constructing a packing of sufficient rate. We remark here that our scheme is closely related to the common randomness generation scheme for hypergraphical sources appearing in [HTKZPW].
Secondly, we give a perfect secret key generation scheme (not necessarily capacity achieving) for generic $3$-uniform hypergraphs. The key building block of this scheme consists of a 2-bit perfect secret key generation scheme for star hypergraphs whose projected graph is a cycle. This is then exploited to show that any star hypergraph generates a perfect secret key of size two times the Hamiltonian packing number of its projected graph. The final scheme then involves packing the original hypergraphs with star hypergraphs whose projected graphs have large Hamiltonian packing numbers. This scheme is shown to be capacity achieving for certain classes of 3-uniform hypergraphs.
II Technical Preliminaries
II-A Multiterminal Source Model
The multiterminal source model of Csiszár and Narayan [CN04] considers a set $\mathcal{M}=\{1,2,\dots,m\}$ of $m\geq 2$ parties. Party $i\in\mathcal{M}$ observes an $n$-length random vector $X_{i}^{n}$ taking values in a finite alphabet $\mathcal{X}_{i}^{n}$. The collection of observations across all parties is denoted by $X_{\mathcal{M}}^{n}=(X_{i}^{n}:i\in\mathcal{M})$. We assume the source is memoryless, meaning the joint distribution factors as $P_{X_{\mathcal{M}}^{n}}(x_{\mathcal{M}}^{n})=\prod_{t=1}^{n}P_{X_{\mathcal{M}}}(x_{\mathcal{M},t})$, where $P_{X_{\mathcal{M}}}$ is a joint probability mass function on $\prod_{i\in\mathcal{M}}\mathcal{X}_{i}$. For any subset $A\subseteq\mathcal{M}$, we define the notation $X_{A}=(X_{i}:i\in A)$.
To generate secret keys, the parties are allowed to (interactively) communicate over a noiseless public channel. The random variable corresponding to the public communication is denoted by $F:=(F_{1},\dots,F_{r})$, where $F_{j}$ denotes the $j$th transmission taking values in some finite set $\mathcal{F}_{j}$. The set of all possible values of the public communication is denoted by $\mathcal{F}=\prod_{i\in r}\mathcal{F}_{r}$. The communication being interactive, the $j$th transmission, sent by say party $i$, is a function of party $i$’s local observation $X_{i}^{n}$ as well as past public messages $F^{j-1}$. Following the public messages each party $i\in\mathcal{M}$ computes a local key $K_{i}=\varphi_{i}(X_{i}^{n},F)$ using a deterministic mapping $\varphi_{i}:\mathcal{X}_{i}^{n}\times\mathcal{F}\to\mathcal{K}$, where the finite set $\mathcal{K}$ is referred to as the key alphabet. [^2] Observe that the random variables $F,K_{i}$ and the sets $\mathcal{F},\mathcal{K}$ implicitly depend on $n$.
Definition 1.
A number $R\geq 0$ is an achievable (strong) secret key (SK) rate if, for every $\epsilon>0$ and for all sufficiently large $n$, there exist a public communication $F$, local keys $K_{i},i\in\mathcal{M}$ and a random variable $K$ (referred to as the secret key) taking values in the alphabet $\mathcal{K}$ such that the following conditions hold:
  1. 1.
    Recoverability: For each party $i\in\mathcal{M}$, the local key $K_{i}$ computed from $(X_{i}^{n},F)$ satisfies $\Pr(K_{i}=K)\geq 1-\epsilon$.
  2. 2.
    Strong Secrecy: The secret key $K$ satisfies $I(K;F)\leq\epsilon$.
  3. 3.
    Almost uniformity: The secret key $K$ satisfies $\frac{1}{n}H(K)\geq\frac{1}{n}\log\mid\mathcal{K}\mid-\epsilon$. [^3] Throughout this paper, logarithms are taken base $2$, and all entropy quantities are measured in bits.
  4. 4.
    Secret Key Rate: The secret key $K$ satisfies $\frac{1}{n}H(K)\geq R-\epsilon$.
The (strong) secret key capacity, denoted by $C_{\mathrm{SK}}(X_{\mathcal{M}})$, is the supremum of all achievable strong SK rates.
Csiszár and Narayan [CN04] evaluated the secret key capacity $C_{\mathrm{SK}}(X_{\mathcal{M}})$ for a generic multiterminal source. More relevant to us is the alternate characterization of $C_{\mathrm{SK}}(X_{\mathcal{M}})$ given by Chan and Zheng in [CZ10]. Let $\mathcal{P}$ denote the set of all partitions of $\mathcal{M}$ in at least two parts. The secret key capacity is then given by
$C_{\mathrm{SK}}(X_{\mathcal{M}})=\min_{\mathcal{P}:|\mathcal{P}|>1}I_{P}(X_{\mathcal{M}}),$ (1)
where
$I_{P}(X_{\mathcal{M}}):=\frac{1}{|\mathcal{P}|-1}\left(\sum_{C\in\mathcal{P}}H(X_{C})-H(X_{\mathcal{M}})\right).$ (2)
Of relevance to us is the partition of $\mathcal{M}$ into singletons, denoted by $\mathcal{S}=\{\{1\},\{2\},\ldots,\{m\}\}$, which we shall henceforth refer to as the singleton partition. Following the convention of [mukherjee], we shall refer to the sources where $\mathcal{S}$ is a minimizer of the right hand side of (1) as Type-$\mathcal{S}$ sources. Therefore, $C_{\mathrm{SK}}(X_{\mathcal{M}})=I_{\mathcal{S}}(X_{\mathcal{M}})$ for Type-$\mathcal{S}$ sources.
While $C_{\mathrm{SK}}(X_{\mathcal{M}})$ is defined under the asymptotic notion of strong secrecy (where error and leakage vanish as $n\to\infty$), there is a stricter non-asymptotic notion known as perfect secrecy. A random variable $K$ is a perfect secret key if it satisfies exact recoverability ($\Pr(K_{1}=\dots=K_{m}=K)=1$), perfect secrecy ($I(K;F)=0$), and perfect uniformity ($H(K)=\log|\mathcal{K}|$) for some fixed blocklength $n$. It is immediate that a perfect secret key satisfies strong secrecy, but not vice versa. Accordingly, the perfect secret key capacity is upper bounded by the usual strong secret key capacity.
We remark that henceforth, the phrase ‘secret key capacity’ will be used to refer to the strong secret key capacity. Furthermore, we shall say that a perfect secret key achieves capacity if it has a rate equal to the secret key capacity. This work shows that for instances of a special class of sources, called the hypergraphical sources, perfect secret keys achieve capacity.
II-B Hypergraphical Source Model
A hypergraphical source is defined using an underlying hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$ whose vertex set is the set of parties $\mathcal{M}$. Associated with each hyperedge $e\in\mathcal{E}$ are $n$ i.i.d. copies of a Bernoulli(1/2) random variable $\xi_{e}$. Furthermore, the collection of random variables $(\xi_{e}:e\in\mathcal{E})$ is mutually independent. The observation of party $i\in\mathcal{M}$ is the collection of variables associated with hyperedges incident to it, i.e., $X_{i}=(\xi_{e}:i\in e)$. Consequently, the joint source is $X_{\mathcal{M}}=(\xi_{e}:e\in\mathcal{E})$.
The combinatorial structure inherent in hypergraphical sources greatly simplifies the expression for the secret key capacity. This is especially evident in the case of the so-called $t$-uniform hypergraphs. A hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$ is said to be $t$-uniform if all hyperedges are of size $t$, i.e., $|e|=t$ for all $e\in\mathcal{E}$. In particular, it is easy to see that the secret key capacity for Type-$\mathcal{S}$ sources on $t$-uniform hypergraphs is given by
$C_{\mathrm{SK}}(X_{\mathcal{M}})=I_{\mathcal{S}}(X_{\mathcal{M}})=\frac{t-1}{m-1}|\mathcal{E}|.$ (3)
A special case of the hypergraphical source is the pairwise independent network (PIN) model [pinorig, nitinrawat], where the underlying hypergraph is a graph. Nitinawarat and Narayan [nitinrawat] has shown that the PIN model admits a linear communication scheme which achieves a linear perfect secret key, via spanning tree packing of the underlying graph. In this work, we develop similar linear perfect secret key generation schemes for $t$-uniform hypergraphical sources, by exploiting the combinatorial structure of the underlying hypergraph.
To proceed, we define some notation relevant to hypergraphical sources. Given a set $A\subseteq\mathcal{M}$, we define the notation $\mathcal{E}_{A}:=\{e\in\mathcal{E}:e\cap A\neq\emptyset\}$. When $A=\{i\}$ (resp., $A=\{i,j\}$) we shall abuse notation and write $\mathcal{E}_{i}$ instead of $\mathcal{E}_{\{i\}}$ (resp., $\mathcal{E}_{i,j}$ instead of $\mathcal{E}_{\{i,j\}}$). For the $n$ random bits $\xi_{e}^{n}$ associated with the hyperedge $e$, we shall use the notation $\xi_{e(j)}$ to denote the $j$th i.i.d. copy of $\xi_{e}$. In fact, whenever necessary, we shall associate $n$ fictitious copies $e(1),e(2),\ldots,e(n)$ with every hyperegde $e\in\mathcal{E}$, and associate $\xi_{e(j)}$ with $e(j)$. Furthermore, we define the set of all $n$ fictitious copies of the hyperedges $\mathcal{E}^{(n)}=\{e(j):e\in\mathcal{E},1\leq j\leq n\}$. Next, given any $\mathcal{E}^{\prime}\subseteq\mathcal{E}$, we define the tuples $\xi_{\mathcal{E}^{\prime}}:=(\xi_{e}:e\in\mathcal{E}^{\prime})$ and $\xi_{\mathcal{E}^{\prime}}^{n}:=(\xi_{e(j)}:e\in\mathcal{E}^{\prime},1\leq j\leq n)$.
Hypergraphical sources in general admit linear communication schemes that consist of every party sending some linear combinations of the bits associated with the hyperedges incident on it. Formally, let $X_{\mathcal{M}}$ be a hypergraphical source defined on the underlying hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$. A public communication $F=(F_{1},F_{2},\ldots,F_{m})$ is said to be linear if for every $i\in\mathcal{M}$, the party $i$ sends the communication $F_{i}=M_{i}\xi_{\mathcal{E}}^{n}$, [^4] All matrix operations and ranks considered in this paper are over $\mathbb{F}_{2}$. where the matrix $M_{i}\in\{0,1\}^{r_{i}\times n|\mathcal{E}|}$ for some integer $r_{i}\geq 0$. Furthermore, the matrices $M_{i}$ must satisfy $M_{i}(r,e(j))=1$ only if $e\in\mathcal{E}_{i}$, to ensure that party $i$’s message involves linear combinations of its local observation only. Observe that the size of this linear communication is $\sum_{i\in\mathcal{M}}r_{i}$ bits. A linear communication is said to achieve perfect omniscience if every party $i$ can compute $\xi_{\mathcal{E}}^{n}$ correctly with probability one using only $X_{i}^{n}$ and $F$.
We next show that any perfect omniscience scheme can be leveraged to generate a perfect secret key. To proceed, we need the following technical lemma.
Lemma 1.
Let $s\in\mathbb{N}$ and let $Y=(Y_{1},Y_{2},\ldots,Y_{s})^{T}\sim\text{unif}\{\{0,1\}^{s}\}$. Let $M\in\{0,1\}^{l\times s}$ be a matrix whose rows we denote as $M_{1},M_{2},\ldots,M_{l}$. For any $1\leq l^{\prime}<l$ if $\text{span}\{M_{1},\ldots,M_{l^{\prime}}\}\cap\text{span}\{M_{l^{\prime}+1},\ldots,M_{l}\}=\{0\}$,5550 here denotes the vector $(\underbrace{0,0,\ldots,0}_{s\text{ times}})$. then $I(M_{1}Y,\ldots,M_{l^{\prime}}Y;M_{l^{\prime}+1}Y,\ldots,M_{l}Y)=0$. Furthermore, $H(M_{l^{\prime}+1}Y,\ldots,M_{l}Y)=\dim(\text{span}\{M_{l^{\prime}+1},\ldots,M_{l}\})$.
Proof:
[…]
Then, by the result stated above, we have H(M_l’+1Y,…,M_lY)=H(BY)=rank(B)=dim(span{M_l’+1,…,M_l}). Also,
$I(M_{1}Y,\ldots,M_{l^{\prime}}Y;M_{l^{\prime}+1}Y,\ldots,M_{l}Y)$ $=I(AY;BY)$
$=H(AY,BY)-H(AY)-H(BY)$
$=H(MY)-H(AY)-H(BY)$
[…]
$=0.$
Lemma 5 is now leveraged to derive lower bounds on the perfect secret key rate achievable through perfect omniscience obtained via linear communication.
Lemma 2.
Let $X_{\mathcal{M}}$ be a hypergraphical source defined on the underlying hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$. For a blocklength $n\in\mathbb{N}$ let there exists a linear communication scheme of $r<n|\mathcal{E}|$ bits achieving perfect omniscience. Then there exists a perfect secret key of rate at least $|\mathcal{E}|-\frac{r}{n}$.
Proof:
[…]
To see that claim holds, firstly observe that $K$ is perfectly recoverable as every party $i\in\mathcal{M}$ has access to $\xi_{\mathcal{E}}^{n}$ thanks to perfect omniscience. Lemma 5 shows that $I(K;F)=0$. Furthermore, as $\{M^{\prime}_{r^{\prime}+1},\ldots,M^{\prime}_{n|\mathcal{E}|}\}$ is linearly independent, Lemma 5 shows that $H(K)=n|\mathcal{E}|-r^{\prime}=\log|\mathcal{K}|$. Thus, the key $K$ is perfectly uniform, and hence $K$ is a perfect secret key. The rate of $K$ is thus $\frac{1}{n}H(K)=|\mathcal{E}|-\frac{r^{\prime}}{n}\geq|\mathcal{E}|-\frac{r}{n}$. ∎
We conclude this section by showing how the notion of packing can be used in tandem with the perfect secret key generation scheme of Lemma 2 to increase the perfect secret key rate available from hypergraphical sources.
A hypergraphical source $X_{\mathcal{M}}$ on the underlying hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$ is said to admit a packing for blocklength $n$, if there exists disjoint subsets $\mathcal{E}_{(1)},\mathcal{E}_{(2)},\ldots,\mathcal{E}_{(s)}\subseteq\mathcal{E}^{(n)}$ satisfying the constraint that for every $1\leq l\leq s$, the set $\mathcal{E}_{(l)}$ contains at most one single fictitious copy of a hyperedge $e\in\mathcal{E}$. The last constraint ensures that for every $1\leq l\leq s$, $\mathcal{H}_{(l)}:=(\mathcal{M},\mathcal{E}_{(l)})$ can be viewed as a hypergraph in its own right. Moreover, we define $X_{i_{(l)}}:=(\xi_{e(j)}:i\in e,e(j)\in\mathcal{E}_{(l)})$, and consequently $X_{\mathcal{M}_{(l)}}:=(X_{i_{(l)}}:i\in\mathcal{M})$ is a hypergraphical source.
Lemma 3.
Let $X_{\mathcal{M}}$ be a hypergraphical source on the underlying hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$. For a blocklength $n$, let $\mathcal{H}$ admits a packing, such that the hypergraphical source $X_{\mathcal{M}_{(l)}},1\leq l\leq s$, can generate a perfect secret key of $k_{l}$ bits using a blocklength of one. Then, $X_{\mathcal{M}}$ can generate a perfect secret key of rate $\frac{\sum_{l=1}^{s}k_{l}}{n}$.
Proof:
Let $K_{l},1\leq l\leq s$, be the perfect secret key of $k_{l}$ bits generated by $X_{\mathcal{M}_{(l)}}$ through public communication $F_{(l)}$. Thus, the overall public communication is $F=(F_{(l)}:1\leq l\leq s)$. Define $K=(K_{l}:1\leq l\leq s)$.
Now, observe that $\xi_{\mathcal{E}_{(1)}},\xi_{\mathcal{E}_{(2)}},\ldots,\xi_{\mathcal{E}_{(s)}}$ are mutually independent as $\mathcal{E}_{(1)},\mathcal{E}_{(2)},\ldots,\mathcal{E}_{(l)}$ are disjoint sets. As a consequence, $K$ is uniformly distributed, $I(K;F)=0$, and $H(K)=\sum_{l=1}^{s}k_{l}$. Hence, $K$ is the required perfect secret key, and its rate is $\frac{\sum_{l=1}^{s}k_{l}}{n}$. ∎
III Secret Key Construction Scheme for complete $t$-Uniform Hypergraphical Sources
In this section, we specialize the general hypergraphical model to the case where $\mathcal{H}$ is a complete $t$-uniform hypergraph, i.e., $\mathcal{E}=\binom{\mathcal{M}}{t}$. We denote the hyperedge variables by $\xi_{i_{1}i_{2}\ldots i_{t}}$ for $\{i_{1}i_{2}\ldots i_{t}\}\in\mathcal{E}$. We assume access to $t$ i.i.d. copies of the source, denoted $X_{\mathcal{M}}^{(1)},\dots,X_{\mathcal{M}}^{(t)}$. These $t$ copies constitute the necessary resources for the structural decomposition.
Definition 2 (Star Model $S_{m,i}$).
Consider a hypergraphical source defined on a complete $t$-uniform hypergraph $\mathcal{H}=(\mathcal{M},\mathcal{E})$. Fix a party $i\in\mathcal{M}$. The Star model $S_{m,i}$ is a pairwise source model defined on the vertex set $\mathcal{M}$ with edge set $\mathcal{E}_{i}=\big\{e^{\prime}\in\mathcal{E}\mid i\in e^{\prime}\big\}$. The random variable associated with an edge $e\in\mathcal{E}_{i}$ is the hyperedge variable $\xi_{e}$ from the original source.
Note that party $i$ observes all edge variables in the projected graph $S_{m,i}$ (since $i$ is a member of every hyperedge). Our construction proceeds in two phases:
  1. 1.
    Global Decomposition: We first define a bijection that partitions the aggregate resources from $t$ copies of the hypergraph into $m$ edge-disjoint Star models.
  2. 2.
    Local Protocol Execution: Once the Star models are defined, each party $i\in\mathcal{M}$ independently executes the secret key generation protocol on the specific Star model $S_{m,i}$ assigned to it.
{restatable}
[Global Source Decomposition]lemmasourcedecomp Given $t$ independent copies of a complete $t$-uniform hypergraphical source $\mathcal{H}=(\mathcal{M},\mathcal{E})$, the aggregate collection of hyperedge copies can be perfectly partitioned into $m$ edge-disjoint Star models $\{S_{m,1},S_{m,2},\dots,S_{m,m}\}$, where $S_{m,i}$ is anchored at party $i$.
1
2
Input: Observations from $t$ independent copies of the source: $X_{\mathcal{M}}^{(1)},\dots,X_{\mathcal{M}}^{(t)}$.
Output: Public communication tuple $F=(F^{(1)},F^{(2)},\dots,F^{(m)})$.
3
Init: Initialize empty communication lists $F^{(i)}\leftarrow\emptyset$ for all $i\in\mathcal{M}$.
4
5foreach hyperedge $e\in\mathcal{E}$ do
6   Fix an ordering of the vertices of $e$ as $(u_{1},u_{2},\ldots,u_{k})$, $k\in\{1,\ldots,t\}$.
7   Assign the $k$th copy of the hyperedge $e^{(k)}$ to the Star model $S_{m,u_{k}}$ anchored at $u_{k}$.
8
9
10foreach party $i\in\mathcal{M}$ do
11   Let $\mathcal{E}_{i}$ be the set of edge variables assigned to $S_{m,i}$.
12   Choose an arbitrary helper party $j\in\mathcal{M}\setminus\{i\}$. Let $\mathcal{E}_{ij}=\{e\in\mathcal{E}\mid i\in e,j\in e\}$.
13
14 foreach $e\in\mathcal{E}_{i}\setminus\mathcal{E}_{ij}$ do
15    $M_{\text{xor}}\leftarrow 0$.
16
[…]
18
19       $M_{\text{xor}}\leftarrow M_{\text{xor}}\oplus\xi_{\{i,j\}\cup A}$.
20
21    $msg\leftarrow\xi_{e}\oplus M_{\text{xor}}$.
22      Append msgmsg to $F^{(i)}$.
23
24  Party $i$ sends $F^{(i)}$.
Algorithm 1 Global Secret Key Construction Protocol
In Algorithm 1, the global execution consists of $m$ independent instances of a Star protocol. For a specific party $i$ acting as the anchor for $S_{m,i}$, the number of messages sent is $\binom{m-1}{t-1}-\binom{m-2}{t-2}$. We define the secret key $K^{(i)}$ generated by this specific instance as the subset of edges in $S_{m,i}$ known to both $i$ and its helper $j$. By the construction of the Star model, this set $K^{(i)}$ corresponds to the variables in $\mathcal{E}_{ij}$, which has cardinality $\binom{m-2}{t-2}$. The subsequent lemmas formally prove that for any such instance, $K^{(i)}$ is perfectly recoverable by all parties and perfectly secret from the public communication.
{restatable}
[Perfect Recoverability]lemmaperfectrec For a fixed anchor $i$, after the execution of Algorithm 1, every party $v\in\mathcal{M}$ can perfectly recover the secret key $K^{(i)}$.
{restatable}
[Perfect Secrecy and Uniformity]lemmaperfectsec The secret key $K^{(i)}$ generated by Algorithm 1 is uniformly distributed and perfectly secret with respect to the public communication $F$. That is, $H(K^{(i)})=\log|\mathcal{K}|$ and $I(K^{(i)};F)=0$.
Theorem 4 (Main Result).
The protocol defined by Algorithm 1 constitutes a capacity-achieving perfect secret key generation scheme for complete $t$-uniform hypergraphical sources.
Proof:
Our protocol begins with the Global Decomposition phase described in Lemma 2. We utilize $t$ copies of the source to construct exactly $m$ edge-disjoint Star models $S_{m,1},\dots,S_{m,m}$, where $S_{m,i}$ is anchored at party $i$. This decomposition is deterministic and exhaustive, consuming all edge resources without collision. We then assume, without loss of generality, that the indices $\{1,\dots,m\}$ assigned to the Star models correspond to the party indices. We execute Algorithm 1 on each of the $m$ Star models in parallel.
For the instance $S_{m,i}$, party $i$ acts as the sender and selects an arbitrary helper $j_{i}\in\mathcal{M}\setminus\{i\}$. From Lemmas 1 and 1, each instance generates a key $K^{(i)}$ that is perfectly recoverable, perfectly secret, and uniform. Since the Star models are edge-disjoint (operating on distinct copies of variables), the execution of the $m$ protocols is mutually independent. Thus, the aggregate properties hold for the combined key.
To calculate the rate, we sum the key bits generated by each model. For a Star model anchored at $i$, the key size is equal to the number of shared edges with helper $j_{i}$, denoted $|\mathcal{E}_{ij}|$. In a complete $t$-uniform hypergraph, for any pair of parties $i,j$, the number of edges containing both is exactly $\binom{m-2}{t-2}$. Thus, $|K^{(i)}|=\binom{m-2}{t-2}$ for all $i$. Since there are $m$ such models, the total number of key bits generated is $m\times\binom{m-2}{t-2}$. The rate is normalized by the $t$ copies of the source utilized in the decomposition $R_{\mathrm{SK}}=\frac{m}{t}\binom{m-2}{t-2}$. Finally, we verify optimality. The secret key capacity for complete $t$-uniform hypergraphs is given as $C_{\mathrm{SK}}(X_{\mathcal{M}})=\frac{t-1}{m-1}\binom{m}{t}=\frac{m}{t}\binom{m-2}{t-2}$. Since $R_{\mathrm{SK}}=C_{\mathrm{SK}}(X_{\mathcal{M}})$, the scheme is capacity achieving. ∎
IV Secret Key Construction Scheme for $3$-uniform Hypergraphs
In this section, we provide a new secret key generation scheme for $3$-uniform hypergraphs (not necessarily complete), and we identify a subclass where our scheme achieves capacity. We start by defining the notion of an induced cycle 33-uniform hypergraph, which is used as a key construct in our scheme.
Definition 3 (Induced cycle 33-uniform hypergraphical source).
A $3$-uniform hypergraphical source of the form $\mathcal{H}_{i}=\{\mathcal{M},\mathcal{E}\}$ is said to be an induced cycle 33-uniform hypergraphical source if: $i\in e\ \forall\ e\in\mathcal{E}$ and the projected graph from ii [^6] For a hypergraph $\mathcal{H}=\{\mathcal{M},\mathcal{E}\}$, the projected graph from $i\in\mathcal{M}$ is the graph $G=\{\mathcal{M}\setminus\{i\},\mathcal{E}_{i}\}$ where $\mathcal{E}_{i}=\{e\setminus\{i\}:e\in\mathcal{E},i\in e\}$. (called the anchor vertex) forms a cycle.
{restatable}
[Secret Key Capacity for the induced cycle 33-uniform hypergraph]propositionprop There exists a scheme for the induced cycle 33-uniform hypergraph that generates $2$ secret key bits, depending on the parity of $m$. Let k=m1k=m-1, then the scheme is given as follows:
  1. 1.
    If $k=3l$ or $k=3l+2$ the anchor $i$ sends the following k2k-2 messages: $e_{1}\oplus e_{2}\oplus e_{3}$, $e_{2}\oplus e_{3}\oplus e_{4}$, … $e_{k-2}\oplus e_{k-1}\oplus e_{k}$.
  2. 2.
    If $k=3l+1$, the anchor $i$ sends the messages $e_{1}\oplus e_{2}\oplus e_{3}$, …, $e_{k-3}\oplus e_{k-2}\oplus e_{k-1}$, and finally the message $e_{k-1}\oplus e_{k}\oplus e_{1}$.
Now, we describe the generic scheme for $3$-uniform hypergraphs.
Definition 4 (Generic Scheme for $3$-uniform hypergraphs).
Consider a 33-uniform hypergraph $\mathcal{H}=\{\mathcal{M},\mathcal{E}\}$.
  1. 1.
    Fix a set of anchor points $A\subseteq\mathcal{M}$.
  2. 2.
    For each anchor $i\in A$, suppose $k_{i}$ copies of the projected graph from $i$ admit a Hamiltonian decomposition of size $p_{i}$.
  3. 3.
    Each edge $e$ is used $t_{e}$ times where $t_{e}=\sum_{i\in A}k_{i}\mathbf{1}\{i\in e\}$, choose $n=\max_{e\in\mathcal{E}}t_{e}$.
  4. 4.
    Consider $n$ independent copies of the original source $\mathcal{H}$ and for each anchor point $i$, consider the $p_{i}$ Hamiltonian decompositions and observe that each decomposition corresponds to an induced cycle hypergraph. We can now run the scheme described in Proposition 3 as a subroutine on each of these.
Observe that since each run of the subroutine gives us $2$ bits of secret key (see Proposition 3), we get a scheme with rate $\frac{2}{n}\sum_{i\in A}p_{i}$. We now identify three different classes of $3$-uniform hypergraphs for which the scheme described in Defintion 4 is capacity achieving.
First we consider the class of complete $3$-uniform hypergraph on $m$ vertices. {restatable}corollarytypea For the complete $3$-uniform hypergraph $K_{m,3}$ on $m$ vertices, applying the generic scheme (see Definition 4) with the anchor set $A=[m]$ achieves capacity.
Next, we consider a particular family of $3$-uniform hypergraphs whose projected graphs admit Hamiltonian decompositions. An example of such a hypergraph are hypergraphs wherein the projected graph from a vertex is a Paley graph [ALSPACH2012113].
{restatable}
corollarytypeb Consider a hypergraphical source $\mathcal{H}=\{\mathcal{M},\mathcal{E}\}$ constructed such that there exists a specific vertex $i\in\mathcal{M}$ where $i\in e$ for all $e\in\mathcal{E}$, such that $k$ independent copies of the projected graph from $i$ admit a Hamiltonian cycle packing of size $p$. Applying the generic scheme with the $i$ as the sole anchor achieves capacity. Further, we show that such a hypergraphical source is Type-S.
Finally, we identify a class of $3$-uniform hypergraphs that we term Hollow 3D Kite hypergraphs. We show that the generic scheme achieves capacity and such a hypergraphical source is Type-S.
{restatable}
corollarytypec Consider a hypergraphical source with $m=4r+1$ vertices. Let the vertex set be partitioned into two sets: $V_{1}=\{1,\dots,2r+1\}$ and $V_{2}=\{2r+2,\dots,4r+1\}$. The edge set is defined as $\mathcal{E}=\left\{\{i,j,k\}:\{i,j\}\subseteq V_{1},k\in V_{2}\right\}$. The generic scheme applied with the anchor set $A=V_{1}$ achieves capacity. Further, such a hypergraphical source is Type-S.
Appendix A Ommitted Proofs
A-A Omitted Proofs of Section III
[…]
Proof:
Note that exactly $t$ vertices participate in every hyperedge $e\in\mathcal{E}$, and there are $t$ independent copies of every hyperedge. Fix an arbitrary ordering of vertices in every hyperedge. For each hyperedge, assign the $k$th copy of hyperedge $e$ to the $k$th vertex. Now consider any vertex $v\in\mathcal{M}$. The vertex $v$ receives exactly one copy of every hyperedge it is a part of, and hence is assigned the edge set $\mathcal{E}_{v}$. This is exactly the star model $S_{m,v}$. Since the choice of vertex was arbitrary, we have $m$ different star models each anchored at a different vertex. ∎
[…]
Proof:
Let $j$ be the helper vertex fixed in Algorithm 1, so the secret key is $K^{(i)}=\mathcal{E}_{ij}$. We analyze the recoverability for three distinct cases covering all parties in $\mathcal{M}$.
  1. 1.
    Case 1: The Anchor Party $i$. Party $i$ is a member of every hyperedge in the Star model $S_{m,i}$. Consequently, $i$ directly observes every variable $\xi_{e}$ for all $e\in\mathcal{E}_{i}$, including the subset $\mathcal{E}_{ij}$. Thus, $i$ possesses $K^{(i)}$ directly from its source observations.
  2. 2.
    Case 2: The Helper Party $j$. The secret key is defined as the set of edges $\mathcal{E}_{ij}=\{e\in\mathcal{E}\mid i\in e,j\in e\}$. By definition, party $j$ is a member of every hyperedge in this set. Therefore, $j$ directly observes every variable $\xi_{e}\in K^{(i)}$. (Note: Although $j$ also observes the public messages $M_{e}$, it does not need them to recover the key; it uses the key and the messages to recover the remaining variables in $\mathcal{E}_{i}\setminus\mathcal{E}_{ij}$, effectively reconstructing the entire Star).
  3. 3.
    Case 3: Any Other Party $k\in\mathcal{M}\setminus\{i,j\}$. We show that $k$ can recover any specific key bit $\xi_{E^{*}}\in K^{(i)}$, where $E^{*}\in\mathcal{E}_{ij}$. If $k\in E^{*}$, then $k$ observes $\xi_{E^{*}}$ directly. If $\notin E^{*}, kk$ must recover $\xi_{E^{*}}$ from the public communication. The algorithm transmits a message $M_{e}$ for the hyperedge $e=(E^{*}\setminus\{j\})\cup\{k\}$. Note that $e\in\mathcal{E}_{i}\setminus\mathcal{E}_{ij}$ because it contains $i$ (from $E^{*}$) and $k$, but explicitly excludes $j$. Since $k\in e$, party $k$ knows variable $\xi_{e}$. The message structure is:
    $M_{e}=\xi_{e}\oplus\bigoplus_{A\subseteq e\setminus\{i\},|A|=t-2}\xi_{i,j,A}.$
    Party $k$ computes the XOR sum. The sum runs over subsets $A$ of $e\setminus\{i\}$. The set $e\setminus\{i\}$ is exactly $(E^{*}\setminus\{i,j\})\cup\{k\}$. If the subset $A$ is formed by excluding $k$, then $A=E^{*}\setminus\{i,j\}$. The corresponding term is $\xi_{i,j,A}=\xi_{E^{*}}$. This is the target variable $k$ wishes to recover. On the other hand, if the subset $A$ includes $k$, then the corresponding edge $\{i,j\}\cup A$ contains $k$. As a member of this hyperedge, $k$ observes the variable directly.
    Thus, the message $M_{e}$ relates the unknown $\xi_{E^{*}}$ to values known by $k$ (the message $M_{e}$, $\xi_{e}$, and all other terms in the sum). Party $k$ can therefore solve for $\xi_{E^{*}}$. Since this holds for all $E^{*}\in\mathcal{E}_{ij}$, $k$ recovers the full secret key $K^{(i)}$. Consequently, having recovered the key, $k$ can also decode all messages $M_{e}$ to recover the entire Star $\mathcal{E}_{i}$, matching the knowledge of $j$.
[…]
Proof:
Let $\mathbf{K}$ denote the vector of random variables corresponding to edges in $\mathcal{E}_{ij}$, and let $\mathbf{U}$ denote the vector of random variables corresponding to edges in $\mathcal{E}_{i}\setminus\mathcal{E}_{ij}$. Since the entries are i.i.d. Be(1/2), the distribution of $\mathbf{K}$ is uniform over $\{0,1\}^{|\mathcal{E}_{ij}|}$. Thus, $H(K^{(i)})=|\mathbf{K}|=\log|\mathcal{K}|$, implying uniformity.
Let $N=|\mathbf{U}|+|\mathbf{K}|=\binom{m}{t}$, be the total dimension. The random variables in $\mathbf{U}$ and $\mathbf{K}$ are i.i.d. Be(1/2) variables, so the joint vector $(\mathbf{U},\mathbf{K})$ is uniformly distributed over $\mathbb{F}_{2}^{N}$. The public communication $F$ consists of a vector of messages $\mathbf{M}$ generated by the rule $\mathbf{M}=\mathbf{U}\oplus G\mathbf{K}$, where $G$ describes the XOR combinations of variables. We analyze the joint distribution of the messages and the key by considering the augmented linear transformation $T:\mathbb{F}_{2}^{N}\to\mathbb{F}_{2}^{N}$:
$\begin{bmatrix}\mathbf{M}\\ \mathbf{K}\end{bmatrix}=\underbrace{\begin{bmatrix}I_{|\mathbf{U}|}&G\\ 0&I_{|\mathbf{K}|}\end{bmatrix}}_{T}\begin{bmatrix}\mathbf{U}\\ \mathbf{K}\end{bmatrix}$.
By construction, $T$ has full rank. Since the input $(\mathbf{U},\mathbf{K})$ is uniform and $T$ is a bijection, the output $(\mathbf{M},\mathbf{K})$ is uniformly distributed over $\mathbb{F}_{2}^{N}$. Hence that $\mathbf{M}$ and $\mathbf{K}$ are independent, i.e., $H(\mathbf{K}\mid\mathbf{M})=H(\mathbf{K})$, yielding $I(\mathbf{K};\mathbf{M})=0$, implying perfect secrecy.
A-B Omitted Proofs of Section IV
[…]
Proof:
Suppose $|\mathcal{M}|=k+1,\ k\geq 4$ (for $k=3$, we simply have a triangle and sending the addition of all messages suffices). Without loss of generality, let the induced cycle source be $\mathcal{H}_{i}=\{[k]\cup\{i\},\mathcal{E}\}$. We introduce the following notation for the edges of the graph: for each $j\in[k-1],e_{j}:=\{i,j,j+1\}$ and $e_{k}=\{i,k,1\}$.
    Case I: $k=3l$ or $k=3l+2$
    The protocol is as follows: the anchor $i$ sends the following k2k-2 messages: $e_{1}\oplus e_{2}\oplus e_{3}$, $e_{2}\oplus e_{3}\oplus e_{4}$, … $e_{k-2}\oplus e_{k-1}\oplus e_{k}$. Once we show recovery, we know that there exists a secret key of at least $2$ bits by Lemma 2.
    To show recovery, first, consider some vertex $j\neq 1$. By virtue of being a member of those edges, vertex $j$ has access to the bits $e_{j-1},e_{j}$. From these, they can recover the bits $e_{j-2},e_{j+1}$ from the communications $e_{j-2}\oplus e_{j-1}\oplus e_{j}$ and $e_{j-1}\oplus e_{j}\oplus e_{j+1}$, respectively. Inductively, they can now recover all the other bits. Now, for vertex $1$, they have access to edges $e_{1}$ and $e_{k}$. By adding the communications $e_{1}\oplus e_{2}\oplus e_{3}$ and $e_{2}\oplus e_{3}\oplus e_{4}$ along with $e_{1}$, they can recover $e_{4}$. In a similar manner, they can recover every bit of the form $e_{3c+1}$.
    If $k=3l$, the last bit vertex $1$ recovers in this sequence is $e_{3l-2}$, which allows them to recover $e_{3l-1}$ from the communication $e_{3l-2}\oplus e_{3l-1}\oplus e_{3l}$. Vertex $1$ now has the same information as vertex $3l$, which we have shown recovers everything, thus, vertex $1$ can recover all the information.
    Similarly, if $k=3l+2$, the last bit recovered in this sequence is $e_{3l+1}$, which allows it to recover all the knowledge of vertex $3l+2$, which suffices again by the same argument as above.
    Case II: $k=3l+1$
    The protocol is as follows: the anchor $i$ sends the messages $e_{1}\oplus e_{2}\oplus e_{3}$, …, $e_{k-3}\oplus e_{k-2}\oplus e_{k-1}$, and finally the message $e_{k-1}\oplus e_{k}\oplus e_{1}$.
    To show recovery, first, consider some vertex $j\notin\{1,k\}$. By virtue of their edges, vertex $j$ has access to $e_{j-1},e_{j}$. From these, they can recover backwards $e_{j-2},\dots,e_{1}$ using the standard communications (e.g., $e_{j-2}\oplus e_{j-1}\oplus e_{j}$). Similarly, they can recover forward $e_{j+1},\dots,e_{k-1}$. Now, possessing both $e_{1}$ and $e_{k-1}$, they can use the last message $e_{k-1}\oplus e_{k}\oplus e_{1}$ to recover $e_{k}$. Thus, they recover all bits.
    For vertex $1$, they have access to edges $e_{1},e_{k}$. From the last message $e_{k-1}\oplus e_{k}\oplus e_{1}$, they immediately recover $e_{k-1}$. As in Case I, by adding communications $e_{1}\oplus e_{2}\oplus e_{3}$ and $e_{2}\oplus e_{3}\oplus e_{4}$ along with $e_{1}$, they can recover $e_{4}$. Continuing this pattern, they recover every bit of the form $e_{3c+1}$. Since $k=3l+1$, we have k3=3l2k-3=3l-2, which is of this form. Thus, vertex $1$ recovers $e_{k-3}$. Now, knowing $e_{k-3}$ and $e_{k-1}$, vertex $1$ uses the message $e_{k-3}\oplus e_{k-2}\oplus e_{k-1}$ to recover $e_{k-2}$. Vertex $1$ now holds $e_{k-2},e_{k-1}$, which recovers the view of party k1k-1 (who we have shown recovers everything).
    For vertex $k$, they have access to $e_{k-1},e_{k}$. Using the last message $e_{k-1}\oplus e_{k}\oplus e_{1}$, they recover $e_{1}$. They now hold $e_{1},e_{k}$, which recovers the view of party $1$. Since party $1$ recovers everything, party $k$ does as well.
[…]
Proof:
We define k=m1k=m-1 and apply the generic scheme as follows:
  1. 1.
    Anchor Selection: We choose the set of anchors to be all vertices $A=[m]$.
  2. 2.
    Projection Analysis: Consider an arbitrary anchor $i\in A$. The projected graph consists of vertices $m]\setminus\{i\}$ and edges $\{e\setminus\{i\}:i\in e\}$. Since the hypergraph is complete, the projected edges connect every pair of vertices in $m]\setminus\{i\}$. Thus, the projected graph is the complete graph $K_{k}$ (where k=m1k=m-1).
We now proceed by analyzing the decomposition and overlap based on the parity of $k$.
Odd Case ($m$ is even)
    Decomposition: Since $k$ is odd, we invoke Walecki’s Theorem [[alspach2008wonderful], Theorem 5.2], which states that the complete undirected graph $K_{k}$ admits a decomposition into $\frac{k-1}{2}$ edge-disjoint Hamiltonian cycles. Thus, we set $p_{i}=\frac{m-2}{2}$ and use a single copy of the projection ($k_{i}=1$).
    Overlap ($n$): Every edge $e\in\mathcal{E}$ contains exactly $3$ vertices. Since $A=[m]$, all three vertices act as anchors for the edge. With $k_{i}=1$, the total usage count for each edge is $t_{e}=3\times 1=3$. We therefore set $n=3$.
Using the rate formula derived in the generic scheme, we have $R=\frac{2\sum_{i\in A}p_{i}}{n}$. Substituting $|A|=m$, $p_{i}=(m-2)/2$, and $n=3$ gives us $R=\frac{2\cdot m\cdot\frac{m-2}{2}}{3}=\frac{m(m-2)}{3}$. This achieves capacity, as the singleton bound for the complete hypergraph is $\frac{2}{m-1}\binom{m}{3}=\frac{m(m-2)}{3}$.
Even Case ($m$ is odd)
For $k\in\{4,6\}$, we provide explicit decompositions in Figures 2 and 2. For other values:
    Decomposition: Since $k$ is even, we consider the complete directed graph $\overrightarrow{K}_{k}$. Tillson’s Theorem [Tillson1980AHD] states that for $k\notin\{4,6\}$, $\overrightarrow{K}_{k}$ admits a decomposition into k1k-1 edge-disjoint directed Hamiltonian cycles. To utilize this, we treat $2$ copies of the undirected projection as a single complete directed graph. Thus, we set $k_{i}=2$ and obtain a decomposition size of $p_{i}=m-2$.
    Overlap ($n$): Every edge contains 3 anchors. Since we use $k_{i}=2$ copies for every anchor, each edge is counted $3\times 2=6$ times across the projections. We therefore set $n=6$.
Using the rate formula, we have $R=\frac{2\sum_{i\in A}p_{i}}{n}$. Substituting $|A|=m$, $p_{i}=m-2$, and $n=6$ gives us $R=\frac{2\cdot m\cdot(m-2)}{6}=\frac{m(m-2)}{3}$. This rate matches the singleton bound capacity of $\frac{m(m-2)}{3}$.
[…]
(a) Copy 1
[…]
(b) Mixed (Copy 1 + Copy 2)
[…]
(c) Copy 2
Figure 1: Hamiltonian cycle packing for $K_{4}$.
[…]
(d) Copy 1
[…]
(e) Copy 1
[…]
(f) Mixed (uses $1\!\to\!2,\,3\!\to\!6$)
[…]
(g) Mixed (uses $4\!\to\!5$)
[…]
(h) Copy 2
Figure 2: Hamiltonian cycle packing for $K_{6}$.
[…]
Proof:
We apply the generic scheme as follows:
  1. 1.
    Anchor Selection: We choose the singleton set of anchors $A=\{i\}$.
  2. 2.
    Projection Analysis: The projected graph at vertex $i$ consists of vertices $\mathcal{M}\setminus\{i\}$ and edges $\{e\setminus\{i\}:e\in\mathcal{E}\}$.
  3. 3.
    Decomposition: We assume that $k$ independent copies of this projected graph admit a Hamiltonian cycle packing of size $p$. Thus, we have a total decomposition size of $p$ distributed over these copies.
  4. 4.
    Overlap ($n$): We set the number of copies of the original source to $n=k$. Since $i$ is in every edge and we consider $k$ copies of the projection, the overlap aligns with the decomposition parameters.
Using the rate formula derived in the generic scheme, we have $R=\frac{2\sum_{i\in A}p_{i}}{n}.$ Substituting the parameters $p_{i}=p$ and $n=k$ gives us $R=\frac{2p}{k}.$ We now verify that this achieves capacity. First, observe that a single Hamiltonian cycle on the projected graph (which has m1m-1 vertices) contains m1m-1 edges. Therefore, the union of $p$ disjoint cycles contains p(m1)p(m-1) edges. Since these edges are distributed over $k$ copies, the number of edges in a single copy of the hypergraph is $|\mathcal{E}|=\frac{p(m-1)}{k}.$ The multi-information for the singleton partition $X_{S}$, denoted $I(X_{S})$, provides an upper bound on the secret key capacity as I(X_S) = 2m-1—E— = 2m-1 ( p(m-1)k ) = 2pk. Since our achievable rate matches the upper bound derived from the singleton partition, the singleton partition must minimize the expression for capacity. This demonstrates that the source is Type S.
[…]
Proof:
  1. 1.
    Anchor Selection: We choose the set of anchors to be $A=V_{1}$.
  2. 2.
    Projection Analysis: Consider an arbitrary anchor $i\in A$. The projected graph consists of vertices $V\setminus\{i\}$ and edges $\{e\setminus\{i\}:i\in e\}$. Since every edge in $\mathcal{E}$ containing $i$ must have its second vertex in $V_{1}\setminus\{i\}$ and its third vertex in $V_{2}$, the projected edges connect every node in $V_{1}\setminus\{i\}$ to every node in $V_{2}$. Observing that $|V_{1}\setminus\{i\}|=2r$ and $|V_{2}|=2r$, the projected graph is the complete bipartite graph $K_{2r,2r}$.
  3. 3.
    Decomposition: It is a known result that the complete bipartite graph $K_{2n,2n}$ admits a decomposition into $n$ Hamiltonian cycles. Thus, for our projected graph $K_{2r,2r}$, we have a Hamiltonian decomposition of size $p_{i}=r$. We set $k_{i}=1$ (a single copy).
  4. 4.
    Overlap ($n$): Every edge $e\in\mathcal{E}$ is of the form $\{u,v,w\}$ with $u,v\in V_{1}$ and $w\in V_{2}$. Since our anchor set is $A=V_{1}$, the anchors present in edge $e$ are exactly $u$ and $v$. Consequently, every edge is used exactly twice, implying $n=2$.
Using the rate formula derived in the generic scheme, we have $R=\frac{2\sum_{i\in A}p_{i}}{n}$. Substituting our parameters $|A|=2r+1$, $p_{i}=r$, and $n=2$ gives us $R=\frac{2\cdot(2r+1)\cdot r}{2}=r(2r+1).$ We now verify that this achieves capacity. The total number of edges in the graph is $|\mathcal{E}|=\binom{2r+1}{2}\times 2r=r(2r+1)(2r)$. The multi-information for the singleton partition $X_{S}$, denoted $I(X_{S})$, provides an upper bound on the secret key capacity as: $I(X_{S})=\frac{2}{m-1}|\mathcal{E}|=\frac{2}{4r}\left(2r^{2}(2r+1)\right)=r(2r+1)$. Since our achievable rate matches the upper bound derived from the singleton partition, the singleton partition minimizes the expression for capacity. Thus, this source is Type S.